Web Cites Research projects a rate of return of 20% on new projects. Management plans to plow back 30% of all earnings into the firm. Earnings this year will be $3 per share, and investors expect a 12% rate of return on stocks facing the same risks as Web Cites.

a.

What is the sustainable growth rate? (Round your answer to 2 decimal places.)

  Sustainable growth rate %  
b.

What is the stock price? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Stock price $    
c.

What is the present value of growth opportunities (PVGO)? (Round your answer to 2 decimal places.)

  PVGO $    
d.

What is the P/E ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  P/E ratio    
e.

What would the price and P/E ratio be if the firm paid out all earnings as dividends? (Round your answer to 2 decimal places.)

  Price $    
  P/E ratio    
Answer #1
  1. What is the sustainable growth rate?

Sustainable growth rate (g) = ROE * Plow back rate

Where,

Return on equity (ROE) = 20%

Plow back rate = 30%

Therefore,

Sustainable growth rate (g) = 20%* 30% = 6%

  1. What is the stock price?

Expected Dividend per Share D1 = Earnings per Share *Dividend Pay-out rate

Where,

Earnings per Share = $3

Dividend Pay-out rate = 1- plow back rate = 1 -30% = 70%

Therefore

Dividend per Share D1 = $3*70% = $2.10 per share

Required rate of return on stock or cost of equity, k = 12%

Stock price P0 = D1/ (k– g) = $2.10 / (12% -6%) = $2.10 /0.06 = $35.00 per share

  1. What is the present value of growth opportunities (PVGO)?

PVGO = Stock Price - (earnings per share/ cost of equity)

=$35.00- ($3/12%) = $35.00 - $25.00 = $10.00

  1. What is the P/E ratio?

P/E ratio = Stock price / Earning per share

= $35.00 /$3 = 11.67           

  1. What would the price and P/E ratio be if the firm paid out all earnings as dividends?

If the firm paid out all earnings as dividends; dividend pay-out ratio = 100% and D1 = $3

And Sustainable growth rate g = 0%

Therefore, Stock Price = P0 = D1/ k = $3.00 / 12% = $25.00 per share       

And P/E ratio = $25.00/$3 = 8.33